### v1

# Stereo Anything: Unifying Stereo Matching with Large-Scale Mixed Data

Abstract
Stereo matching has been a pivotal component in 3D vision, aiming to find corresponding points between pairs of stereo images to recover depth information. In this work, we introduce StereoAnything, a highly practical solution for robust stereo matching. Rather than focusing on a specialized model, our goal is to develop a versatile foundational model capable of handling stereo images across diverse environments. To this end, we scale up the dataset by collecting labeled stereo images and generating synthetic stereo pairs from unlabeled monocular images. To further enrich the model’s ability to generalize across different conditions, we introduce a novel synthetic dataset that complements existing data by adding variability in baselines, camera angles, and scene types. We extensively evaluate the zero-shot capabilities of our model on five public datasets, showcasing its impressive ability to generalize to new, unseen data (Figure 1). Code will be available at https://github.com/XiandaGuo/OpenStereo.
Figure 1: StereoAnything demonstrates impressive generalization capabilities in different unseen scenes.
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding authors: long.chen@ia.ac.cn
1 Introduction
Computer vision is currently undergoing a revolution due to the development of foundation models in object recognition [43], image segmentation [21], and monocular depth estimation [31, 49, 19, 50], which demonstrate strong zero- and few-shot performance across various downstream tasks. Stereo matching is fundamental for enabling depth perception and 3D reconstruction of observed scenes, playing a critical role in applications such as robotics [55], autonomous driving [16, 15], and augmented reality [47]. However, the exploration of foundation models in stereo matching remains limited due to the extreme difficulty of obtaining accurate disparity ground truth (GT) data. Although numerous stereo datasets [27, 11, 29, 48, 18] have been published, fully leveraging these datasets for training is challenging. Moreover, these annotated datasets are insufficient to train an ideal foundation model even when combined.
Stereo-from-mono [44] is a pioneer study that aimed to generate stereo image pairs and disparity maps directly from monocular images to overcome those challenges. However, this approach resulted in the creation of only 500,000 data samples, which is relatively limited considering the scale required to train robust foundation models effectively. While this effort represents an important step towards reducing the dependency on expensive stereo data collection, the generated dataset is still insufficient for building large-scale models capable of generalizing well to diverse real-world conditions.
In this work, we introduce StereoAnything, a foundation model for stereo matching aimed at providing high-quality disparity estimation for any pair of rectified stereo images, regardless of scene complexity or environmental conditions. Our primary focus is on creating a highly generalized and scalable solution capable of handling diverse scenarios. To achieve this ambitious goal, we adopt a strategy centered on significantly scaling up the training dataset, ensuring that the model is exposed to an extensive variety of scenes. To fully leverage the existing stereo datasets, we utilize numerous publicly available annotated stereo datasets as a core part of our training data, as well as create a new synthetic dataset, StereoCarla, to further increase the quality, quantity and variety of training data available. Furthermore, inspired by the successes of the stereo-from-mono [44] approach and the new monocular depth foundation models [31, 49, 19, 50], we supplement our training with a large volume of synthetic stereo data generated from monocular images. By combining both traditional stereo datasets and newly generated monocular-based stereo pairs, we create a diverse and comprehensive training set that allows StereoAnything to excel in various environments, ensuring robust and accurate depth predictions. This integration of both real and synthetic data helps us overcome limitations related to data scarcity, ultimately pushing the boundaries of stereo-matching capabilities and improving generalization across domains.
Our contributions can be summarized as follows:
  • We first emphasize the importance of scaling up labeled stereo datasets, through an exhaustive study on how different synthetic datasets affect the performance of trained stereo models.
  • Then, we build a new synthetic dataset to better generalize across different scenarios and enhance the performance, called StereoCarla. Compared to existing datasets, it features unique view angles and baselines.
  • We further scale up the training data by effectively incorporating both synthetic stereo data and diverse unlabeled monocular images for training stereo networks.
  • Our final dataset allows for training a stereo model to exhibit the strongest zero-shot capability among all of the existing networks.
2 Related Work
Stereo matching has been extensively studied in the last decade. We identify and review two main bodies of literature, concerning in-domain and cross-domain stereo matching, as well as review techniques for multi-dataset training.
2.1 In-domain Stereo Matching
Early methods depended on hand-crafted features and optimization-based matching [33, 17]. With the rise of deep learning, the CNN-Based model [20, 6, 14, 54] has revolutionized stereo matching by learning feature representations and matching costs directly from data, leading to significant improvements in accuracy and robustness. GCNet [20] firstly constructs concatenation-based 4D cost volumes and utilizes 3D CNNs for cost aggregation. Based on GCNet, PSMNet [6] enhances feature extraction using a spatial pyramid pooling module and employs a stacked hourglass architecture for 3D cost aggregation. GANet [53] accelerates the network by replacing the 3D CNN with a semi-global aggregation layer and a local guided aggregation layer. LEAStereo [9] employs neural architecture search to optimize stereo-matching networks. Recently, there has been a growing interest in video stereo matching [58, 18, 52], particularly in leveraging temporal information to achieve consistent stereo matching. Despite their success, these approaches often exhibit domain-specific biases, limiting their performance on unseen datasets.
Dataset Published Indoor Outdoor Dynamic Video Dense Accuracy Diversity Annotation Baseline FL Range Ave./Med. # images
Sintel [3] ECCV2012 High Medium Synthetic 0.1m - 0-972 66.5/25 1,064
SceneFlow [27] CVPR2016 High High Synthetic 0.54m - 0-10501 53.9/36 35,454
FallingThings [40] CVPRW2018 High Low Synthetic 6cm 768.2px 7-461 35.2/34 61,500
Argoverse [7] CVPR2019 Low Low LiDAR - - 0-256 69.1/59 5,530
VirtualKITTI2 [4] ArXiv2020 High Mid Synthetic 0.53m 725px 0-411 30.1/25 21,260
Tartanair [42] IROS2020 High High Synthetic - - 0-499 21.0/13 306,637
InStereo2K [2] SCIS2020 Low Low LiDAR 10cm(5cm) 8mm 0-328 78.4/74 2,010
UnrealStereo4K [38] CVPR2021 High High Synthetic 0.2m(0.5m) - 0-1515 175.3/135 8,200
CREStereo [23] CVPR2022 High High Synthetic - - 0-2048 15.2/8 200,000
Spring [28] CVPR2023 High Low Synthetic 6.5cm - 0-554 38.1/19 5,000
DynamicStereo [18] CVPR2023 High Low Synthetic 4cm-30cm - 3-35,097 62.7/48 144,900
StereoCarla (Ours) - High High Synthetic 10cm,54cm,100cm, 200cm,300cm 10m 0-2644 75.7/31 552,050
KITTI12 [11] CVPR2012 Medium Low LiDAR 0.54m 720px 4-232 40.1/38 194
Middlebury [34] GCPR2014 High Low LiDAR 140-400mm 1100px-3600px 15-323 72.5/63 15
KITTI15 [29] CVPR2015 Medium Low LiDAR 0.54m 520px 4-230 35.2/33 200
ETH3D [35] CVPR2017 High Low LiDAR 59.5mm-60.4mm 529px-712px 0-62 13.7/10 27
DrivingStereo [48] CVPR2019 Low High LiDAR 0.54m 1003 4-128 31.0/26 7,751
Table 1: Datasets used in our work. Top: Our training sets. Bottom: Our test sets. FL refers to focal length. Ave./Med. refers to the average/median of disparity. Range refers to disparity range.
2.2 Cross-domain Stereo Matching
Domain-invariant representations are crucial for cross-domain stereo matching. Tonioni et al. [37] propose an unsupervised adaptation method that can fine-tune a stereo-matching model without any ground-truth information. DSMNet [54] incorporates a trainable non-local graph filter and domain normalization layer to extract robust structural features. CFNet [36] addresses the large domain difference by introducing a cascade and fused cost volume. FCStereo [56] proposes contrastive learning to maintain feature consistency between matching pixels, which is vital for the generalization of stereo matching. To reduce domain shifts, Chang et al. [8] presents a hierarchical visual transformation network that focuses on learning shortcut-invariant representation from synthetic data. RAFT-Stereo [24] and IGEV [46] introduce multi-level ConvGRUs to update the disparity map and achieve strong cross-dataset generalization iteratively. MS-Net [5], GraftNet [25], and FormerStereo [59] obtain domain-invariant features by leveraging external priors and employing robust feature extractors. The primary difference is that MS-Net [5] and GraftNet [25] focus on CNN backbones, while FormerStereo [59] employs a ViT-based foundational model. NMRF-Stereo [13] presents a neural MRF model, which exhibits strong cross-domain generalization and can recover sharp edges. Zhang et al. [57] explores fine-tuning stereo-matching networks in a manner that preserves their robustness to unseen domains and harnesses the discrepancies between ground truth data and pseudo labels to extract valuable insights for the fine-tuning process.
We refer the reader to recent surveys [22, 30, 39] for further details on both in-domain and cross-domain stereo.
2.3 Multi-Dataset Training
To tackle the challenge of generalization, some researchers have increasingly explored the use of multi-dataset training strategies. By combining multiple datasets, models can be trained to capture diverse scene characteristics better, making them more robust. Although well-established in fields like object recognition [43], segmentation [21], and monocular depth estimation [31, 49, 19, 50], the use of this strategy in stereo matching remains relatively underexplored.
In this work, we introduce StereoAnything, a stereo-matching model that systematically leverages multi-dataset learning. Inspired by the successes in related domains, we aim to enhance the generalization capabilities of stereo-matching models by training on a diverse collection of datasets. By doing so, we seek to develop a model that performs consistently across any conditions.
3 Existing Datasets
We start by revising existing labeled stereo datasets, and unlabeled monocular datasets. Table 1 summarizes the existing labeled datasets involved in our work, dividing them into training and testing sets, along with their properties.
3.1 Training datasets.
3.1.1 Labeled stereo datasets.
To fully leverage the existing stereo datasets, we utilize numerous publicly available annotated stereo datasets as a core part of our training data.
Sintel [3], sourced from computer-generated films and providing dense ground-truth labels, is separated by training datasets (1,064 stereo pairs) and tests (564 stereo pairs).
SceneFlow [27] is a synthetic dataset comprising indoor and outdoor scenes. It is one of the most commonly used datasets in stereo matching, providing valuable ground truth for pre-train stereo-matching algorithms.
CREStereo [23] focuses on challenging scenarios in real-world scenes by incorporating various types of lighting conditions with random colors and luminance at different positions.
FallingThings [40] is a large-scale synthetic dataset designed to advance research in 3D object detection and pose estimation, also providing annotated stereo images.
InStereo2K [2] introduces a dataset of over 2,000 real-world, indoor stereo image pairs with high-quality, semi-dense ground-truth disparities.
Spring [28] provides 6,000 high-resolution image pairs, each with a 2.1 Mpx resolution, spanning 47 diverse scenes.
UnrealStereo4K [38] is a high-resolution, synthetic stereo dataset, including 8Mpx stereo pairs.
Argoverse [7] is a dataset for 3D tracking and motion forecasting. It contains 5,530 forward-facing stereo images for training and 1,094 for testing with a resolution of $2056\times 2464$, collected in real driving environments.
VirtualKITTI2 [4] employs advancements in Unity’s lighting and post-processing techniques to generate a dataset with minimal differences between generated and real images from KITTI. It includes 21,260 stereo pairs.
Tartanair [42] is a large-scale synthetic dataset specifically designed to advance the development of Visual SLAM. It offers a diverse range of simulated environments with varying lighting conditions, weather effects, and dynamic objects, which includes 306,637 stereo pairs.
DynamicStereo [18] distinguishes among others for its extended sequences and the presence of non-rigid objects such as animals and humans. The dataset comprises 484/20/20 sequences for training/validation/testing.
3.1.2 Unlabeled monocular datasets.
Dataset Indoor Outdoor # Images
BDD100K [51] 0.1M
Places365 [60] 2.1M
Google Landmarks [45] 4.97M
ImageNet-21K [32] 13.1M
Table 2: Unlabeled monocular images. M refers to million.
(a) Left
(b) Right (10cm)
(c) Right (54cm)
(d) Right (100cm)
(e) Right (200cm)
(f) Right (300cm)
(g) Depth
(h) Disparity (10)
(i) Disparity (54)
(j) Disparity (100)
(k) Disparity (200)
(l) Disparity (300)
(m) Normal
(n) Roll $5^{\circ}$
(o) Roll $15^{\circ}$
(p) Roll $30^{\circ}$
(q) Height $20m$
(r) Pitch $-30^{\circ}$
Figure 2: Overall of StereoCarla datasets. The first row illustrates the left-eye image (1st column) and the right-eye image at varying baselines (2nd - 6th columns). The second row showcases the depth map (1st column) and corresponding disparity maps (2nd - 6th columns). The third rows depict left images from varied horizontal viewing angles and elevated viewpoints.
To enhance the robustness of the stereo-matching model, we incorporate a large-scale collection of 20.27 million unlabeled monocular images sourced from four public datasets. A comprehensive overview of these datasets is provided in Table 2. These will be used to generate further, synthetic stereo pairs during training.
3.2 Testing datasets.
To evaluate the zero-shot generalization performance of the stereo-matching model, we select five datasets known for their diversity and accurate ground truth annotations.
KITTI 2012 [11] and KITTI 2015 [29] datasets are widely used benchmarks in stereo matching with sparse LiDAR ground truth disparities, featuring real-world images captured from various driving scenarios. KITTI 2012 [11] includes 194 training and 195 testing pairs, while KITTI 2015 [29] has 200 training and 200 testing pairs.
Middlebury [34] focus on indoor environments, offering 15/15 high-resolution stereo pairs for training and testing annotated with semi-dense ground-truth labels obtained with a structured light system.
ETH3D [35] has 27/20 grayscale stereo image pairs for training and testing, collected from both indoor and outdoor scenes, annotated with semi-dense ground-truth labels obtained with a Faro Focus X 330 laser scanner.
DrivingStereo [48] is a large-scale stereo dataset which includes over 180K images covering a diverse set of real-world driving scenarios. Sparse ground-truth annotations are obtained from LiDAR scans, being post-processed through a deep neural network.
4 StereoCarla Datasets
To expand the diversity and quantity of existing stereo-matching datasets, we utilized the CARLA [10] simulator to collect new synthetic stereo data. CARLA [10], a widely used open-source simulator for autonomous driving research, enabled us to create a range of realistic virtual environments, which significantly increased the flexibility of our data collection process. Compared to previous stereo datasets, our approach offers more varied settings, providing different baselines and novel camera configurations that enhance the richness of stereo data. And Figure 2 illustrates some examples of the newly collected dataset, showcasing the diverse baselines, camera angles, and scene variations. Below, we detail the major design considerations:
Multiple Baselines. We collected data with baseline distances set at 10 cm, 54 cm, 100 cm, 200 cm, and 300 cm, which are much broader in range compared to existing datasets. These baseline variations allow the model to better generalize to scenarios where the distance between the two cameras can differ greatly, such as switching between different hardware configurations in real-world applications.
Horizontal Viewing Angles. We collected data at four different horizontal viewing angles. Specifically, we captured images at a direct horizontal view (0 degrees) and slight and moderate rotations of 5 degrees, 15 degrees, and 30 degrees. These different orientations provide more robust data, as they simulate variations that occur naturally when a stereo camera system is in motion or when changes in scene viewpoint occur. This kind of angular variation is crucial for enhancing the robustness of a model, making it more resilient to real-world shifts in perspective.
Elevated Viewpoints. Furthermore, to simulate a diverse set of environments and perspectives, we positioned the stereo camera pair at an elevated height of 10 meters above the vehicle, capturing the scenes from both a horizontal view and at a 30-degree downward tilt. This unique perspective provides more information about the overall layout of the scene, which can be particularly useful in understanding both near-ground details and broader contextual elements that are missing in most existing datasets.
High Resolution. Each scene in our dataset consists of left and right images, with a resolution of $1600\times 900$ pixels, along with corresponding dense disparity maps, ensuring comprehensive ground-truth information for both training and evaluating state-of-the-art stereo models.
By incorporating the above elements, our dataset aims to address key limitations in existing stereo datasets, particularly the lack of varied viewpoints and baseline distances, thereby providing a more robust foundation for developing and testing stereo-matching algorithms.
5 StereoAnything
We now introduce our framework, StereoAnything, designed for training a robust stereo network by using large-scale mixed data.
Figure 3: Stereo Matching Pipeline. Deep stereo models usually perform feature extraction, cost construction, cost aggregation, and disparity regression.
5.1 Pipeline of Stereo Matching
As shown in Figure 3, stereo matching based on deep learning involves estimating disparity from a pair of rectified stereo images. It mainly consists of four essential components [6, 13, 39]: feature extraction, cost construction, cost aggregation, and disparity regression.
5.2 Learning Labeled Stereo Images
In this section, we describe our approach to enhancing stereo model generalization by leveraging supervised stereo data. It is important to note that monocular depth models like MiDaS [31], DepthAnytingV1 [49], and DepthAnytingV2 [50] normalize the depth values to a range of 0 to 1 on each depth map. However, this normalization is clearly unsuitable for stereo matching, as stereo matching has inherent scale information that would be lost during the process.
Our method begins by training the stereo model on a single dataset and subsequently evaluating its generalization performance across test datasets. To quantify the generalization ability, we calculate the mean metric across these five datasets, producing a ranking of datasets that allows us to assess their impact on cross-dataset performance:
$P_{r_{1}}=f_{eval}(M,D_{1})$ (1)
Next, we rank all datasets based on evaluation metrics:
$\{D_{r_{1}},D_{r_{2}},\dots,D_{r_{n}}\},\\ \text{such that}\quad P_{r_{1}}\geq P_{r_{2}}\geq\dots\geq P_{r_{n}}$ (2)
The datasets are ranked from highest to lowest performance.
The mixing process starts with the top-ranked dataset $\mathcal{D}_{r_{1}}$, and we sequentially add other datasets to the mixed training set based on their ranking. This approach ensures that the training process incorporates data that is most beneficial to improving generalization performance, ultimately yielding a model that is more robust when applied to various unseen stereo datasets:
${MIX}_{k}=\bigcup_{i=1}^{k}D_{r_{i}},\quad\text{where }k\in\{1,2,\dots,n\}$ (3)
By employing this incremental combination strategy, we aim to exploit the diversity of multiple datasets while ensuring that the highest-quality data is prioritized during training.
5.3 Learning Stereo from Single Images
This process is similar to work [44]. Given a single color image $I$, a pre-trained monocular depth estimation model $g$ is used to estimate the depth $Depth$ of image $I$:
$Depth=g(I)$ (4)
Next, we need to convert the estimated depth $Z$ into a disparity map $\tilde{D}$. For simulating realistic stereo pairs with varied baselines and focal lengths, the disparity $\tilde{D}$ is calculated as follows:
$\tilde{Disp}=\frac{s\times Depth_{\max}}{Depth}$ (5)
where $s$ is a scaling factor randomly sampled from a uniform distribution $[disp_{\min},disp_{\max}]$, ensuring that the disparities lie within a plausible range.
The goal is to synthesize a new ‘right’ image $\tilde{I}_{r}$ from the original image $I_{l}$ and the predicted disparity $\tilde{Disp}$. Unlike backward warping, forward warping is applied here to generate $\tilde{I}_{r}$, by translating each pixel $i$ in $I_{l}$ by $\tilde{Disp}_{i}$ to the left. This approach requires interpolation for filling in missing values and involves challenges such as occlusion handling and pixel collisions.
Finally, we need to handle occlusions and collisions. During forward warping, certain regions in the right image $\tilde{I}_{r}$ may remain unfilled due to occlusion, resulting in unnatural gaps. To mitigate this, we use textures from a randomly selected image $I_{b}$ in our dataset to fill in the missing pixels. We perform color transfer between $I_{l}$ and $I_{b}$ to match their visual properties before filling.
6 Experiment
We now evaluate the performance achieved by a state-of-the-art stereo model trained within StereoAnything.
6.1 Implementation Details
Our experiments are conducted based on OpenStereo [15]. We adopt NMRF-Stereo [13] as our baseline model due to its strong performance and speed in stereo matching. SwinTransformer [26] is adopted for feature extraction. AdamW optimizer is used for training. For pretraining on SceneFlow [27], we utilize OneCycleLR scheduling with a maximum LR of 0.001, while fine-tuning on other datasets uses OneCycleLR with a maximum LR of 0.0005. The batch size is consistently set to 16 across all experiments. For fine-tuning on different single training sets, we train the model for a total of 31,250 iterations. When fine-tuning on mixed datasets, we extend the training to 81,000 iterations to allow the model to effectively learn from the combined data. For fine-tuning on pseudo stereo datasets, we train the model for one epoch, due to the large amount of data. Data augmentation techniques, including color jitter and random erasing, are applied to enhance the robustness of the model.
6.2 Test datasets and metrics.
For all experiments, we evaluate our method on five benchmark datasets: KITTI2012 [11] (194 training image pairs), KITTI2015 [29] (200 training image pairs), Middlebury [34] (15 image pairs), ETH3D [35] (27 image pairs), and DrivingStereo [48] (7,751 image pairs). Middlebury [34] is tested on half-resolution. Following recent works [54, 46], we use the D1-all metric for KITTI2012 [11], KITTI2015 [29], and DrivingStereo [48], Bad 2.0 for Middlebury [34], and Bad 1.0 for ETH3D [35]. The D1-all computes the percentage of pixels with an absolute disparity error larger than 3.0 pixels. Bad 1.0/Bad 2.0 metric measures the percentage of pixels with a disparity error greater than 1.0/2.0 pixel.
Method K12 K15 Midd E3D DR Mean
PSMNet [6] 30.51 32.15 33.53 18.02 36.19 30.08
CFNet [36] 13.64 12.09 23.91 7.67 27.26 16.91
GwcNet [14] 23.05 25.19 29.87 14.54 35.40 25.61
COEX [1] 12.08 11.01 25.17 11.43 24.17 16.77
FADNet++ [41] 11.31 13.23 24.07 22.48 20.50 18.32
Cascade [12] 11.86 12.06 27.39 11.62 27.65 18.12
LightStereo-L 6.41 6.40 17.51 11.33 21.74 12.68
IGEV [46] 4.88 5.16 8.47 3.53 6.2 5.67
StereoBase [15] 4.85 5.35 9.76 3.12 11.84 6.98
NMRFStereo [13] 4.20 5.10 7.50 3.80 11.92 6.50
NMRFStereo [13] 8.67 7.46 16.36 23.46 34.58 18.11
StereoAnything 4.29 4.31 6.96 1.84 7.64 5.01
Table 3: Comparison with SOTA methods on KITTI12 [11], KITTI15 [29], Middlebury [34], ETH3D [35], and DrivingStereo [48]–lower is better. NMRFStereo refers to NMRFStereo-SwinT. Bold: Best.
Method K12 K15 Midd E3D DR Mean
LightStereo-L 6.41 6.40 17.51 11.33 21.74 12.68
LightStereo-L 4.29 5.82 9.5 5.10 9.44 6.83
NMRFStereo [13] 4.20 5.10 7.50 3.80 11.92 6.50
NMRFStereo [13] 3.50 4.52 6.43 2.35 9.4 5.24
NMRFStereo [13] 8.67 7.46 16.36 23.46 34.58 18.11
NMRFStereo [${}^{*}\dagger$ 4.29 4.31 6.96 1.84 7.64 5.01
Table 4: Ablation study on different models. NMRFStereo refers to NMRFStereo-SwinT. refers to our training strategy.
6.3 Comparison with other SOTA methods
As shown in Table 3, we compare our best-performing model with several SOTA stereo-matching methods. StereoAnything achieves the lowest error rates on the KITTI 2012 [11], KITTI 2015 [29], and ETH3D [35] datasets, outperforming previous methods. In Table 4, we further present an ablation study that highlights the significant impact of our proposed training strategy, which applied to different stereo backbones yields significant improvement in performance across all datasets. Our training approach consistently delivers significant performance improvements across all datasets, showcasing its versatility and effectiveness. The most striking improvement is seen in the NMRFStereo-SwinT [13] variant shows particularly significant improvement, with the mean error reduced from 18.11 to 5.01. This notable gain is attributed to the increased data scale, which allows the SwinTransformer [26] backbone to unlock its full potential.
6.4 Ablation Studies
In this section, we are going to ablate our StereoAnything framework, showing in detail how different datasets – labeled and pseudo-labeled – combinations impact the final results.
Mix KITTI12 [11] KITTI15 [29] Middlebury [34] ETH3D [35] DrivingStereo [48] Mean Rank
D1 EPE Rank D1 EPE Rank Bad2.0 Rank Bad1.0 Rank D1 EPE Rank
SF [27] 8.67 1.67 - 7.46 1.52 - 16.36 - 23.46 - 34.58 10.53 - 18.11 -
MIX 1 4.65 1.19 9 6.01 1.40 9 10.63 9 12.02 8 9.31 2.17 2 8.52 9
MIX 2 4.11 1.04 7 5.19 1.25 7 8.83 8 10.52 7 8.93 2.19 1 7.52 6
MIX 3 4.46 1.11 8 5.33 1.26 8 7.64 6 7.81 5 33.91 6.40 9 11.83 8
MIX 4 3.79 0.91 5 4.38 1.04 5 7.92 7 10.37 6 24.01 4.57 8 10.09 7
MIX 5 3.73 0.90 3 4.81 1.08 6 7.02 4 12.20 9 22.50 4.15 7 10.05 5
MIX 6 3.75 0.79 4 4.29 1.02 4 6.90 3 6.99 3 11.81 2.60 4 6.74 4
MIX 7 3.64 0.78 2 4.28 1.01 3 6.60 1 7.50 4 9.54 2.20 3 6.31 1
MIX 8 3.81 0.79 6 4.21 1.01 2 6.76 2 4.02 1 15.09 4.82 6 6.78 3
MIX 9 3.55 0.77 1 3.70 0.95 1 7.08 5 6.26 2 11.87 3.22 5 6.49 2
Table 5: Cross-domain evaluation on different combinations training datasets. The different MIX setups are explained in Table 7.
Dataset K12 K15 Midd E3D DS Mean Rank
SF [27] 8.67 7.46 16.36 23.46 34.58 18.11 -
[…]
Table 6: Cross-domain evaluation when fine-tuning on different single training sets. Relative to baseline (top row), green/red indicates performance improvement/decline. Bold: Best.
6.4.1 Training on single labeled stereo datasets.
As shown in Table 6, fine-tuning on synthetic datasets such as Sintel [3], FallingThings [40] and Tartanair [42] generally leads to improved performance over the baseline. However, fine-tuning on certain datasets like Argoverse [7] and UnrealStereo4K [38] results in significantly degraded performance, with mean metric increasing to 50.41% and 39.58%, respectively. This suggests that these datasets may not provide sufficient diversity or introduce biases hindering the model’s generalization of different scenarios. In contrast, our proposed dataset, StereoCarla, demonstrates superior generalization capabilities. Fine-tuning on StereoCarla yields the lowest mean metric of 8.52%. Notably, it performs excellently on DrivingStereo [48] with a D1-all metric of 9.31%, highlighting its effectiveness in capturing real-world driving scenarios. In summary, the choice of fine-tuning dataset plays a crucial role in the generalization ability of stereo-matching models. Datasets that closely mimic the diversity and complexity of real-world environments, such as our StereoCarla, significantly enhance performance across multiple benchmarks.
6.4.2 Training on mixed labeled stereo datasets.
In the mixed experiments, we begin with our StereoCarla dataset, as it achieved the best performance in the previous experiment. Then, similar to [31], we sequentially add other datasets to the mixed dataset according to the ranking results in Table 6.
Table 5 and Table 7 show the detailed performance results across these benchmarks for each mix configuration. Notably, the initial combinations (MIX 1 to MIX 3) bring significant improvements in D1 metrics, particularly on KITTI benchmarks. This improvement demonstrates the effectiveness of our StereoCarla dataset as a robust starting point, providing a strong foundation for generalization. As we progress from MIX 4 to MIX 7, incorporating more datasets leads to further performance gains, with MIX 7 yielding the lowest mean error across all tested datasets, achieving a mean D1 error of 6.31, which is the best performance among all configurations. Interestingly, MIX 9, which includes the maximum number of datasets, shows slightly diminished performance compared to MIX 7 in terms of mean error. This suggests that simply adding more datasets may introduce variations that do not always translate into higher accuracy across all benchmarks. These results underscore the importance of carefully selecting and balancing the training datasets to optimize model generalization for diverse real-world scenarios.
Mix Carla DY CRE FT I2K Tar ST SP VK2
MIX 1
MIX 2
MIX 3
MIX 4
MIX 5
MIX 6
MIX 7
MIX 8
MIX 9
Table 7: Combinations of labeled stereo datasets.
6.4.3 Training on pseudo stereo datasets.
As shown in Table 8, we first validate the impact of using different monocular depth estimation models to generate pseudo stereo data on the BDD100k dataset. Fine-tuning on the stereo datasets synthesized using MiDas [31], DepthAnytingV1 [49] (DAv1), and DepthAnytingV2 [50] (DAv2) leads to competitive results, demonstrating the effectiveness of using monocular depth models for pseudo stereo data generation. Specifically, the synthesized data from DAv2 [50] outperforms other monocular methods across all benchmarks.
Figure 4: Qualitative results on five unseen datasets. NMRFStereo refers to NMRF-Stereo-SwinT [13].
Then, we evaluate the impact of using different monocular datasets on the performance of the pseudo stereo data generation, based on the DAv2 [50] model. As shown in Table 9, the Google Landmarks [45] achieves the best performance overall, with a mean metric of 5.91. The slightly worse performance of ImageNet21K [32] (21K) compared to the GL [45] dataset can be attributed to the lower resolution of ImageNet21K images, which is smaller than our model’s input size. During the process, we resized the ImageNet21K [32] images to fit the input size of our model before running the monocular model to estimate depth, followed by generating the pseudo stereo pairs. This resizing process might have led to a loss of fine details, thereby affecting the quality of the generated depth and resulting in slightly lower performance compared to GL [45].
Method K12 K15 Midd E3D DS Mean
MiDas [31] 4.21 4.23 12.04 4.70 9.22 6.88
DAv1 [49] 4.01 3.98 12.08 4.48 8.81 6.67
DAv2 [50] 3.69 3.87 11.75 3.92 9.55 6.56
Table 8: Ablation study of different monocular depth estimation methods.
Dataset K12 K15 Midd E3D DS Mean
SF [27] 8.67 7.46 16.36 23.46 34.58 18.11
BDD [51] 3.69 3.87 11.75 3.92 9.55 6.56
PL [60] 3.78 3.91 10.41 3.65 9.19 6.19
21K [32] 4.10 4.43 10.44 4.25 9.08 6.46
GL [45] 3.86 3.98 9.48 3.26 8.97 5.91
Table 9: Ablation study of using pseudo-stereo generated from different datasets.
Dataset K12 K15 Midd E3D DS Mean
SF [27] 8.67 7.46 16.36 23.46 34.58 18.11
MIX7 3.64 4.28 6.60 7.50 9.54 6.31
GL [45] 3.86 3.98 9.48 3.26 8.97 5.91
M7+GL 4.29 4.31 6.96 1.84 7.64 5.01
Table 10: Cross-domain evaluation combines mix labeled stereo and pseudo-stereo generated from different datasets.
6.4.4 Training on labeled and pseudo stereo datasets.
As shown in Table 10, combining the MIX7 labeled data with pseudo-stereo data from the GoogleLandmarks [45] dataset (denoted as M7+GL-62k) results in further performance gains, achieving the lowest mean metric of 5.01. This combination benefits from both the diversity of labeled stereo data and the robust, diverse depth signals from GoogleLandmarks [45]. These results suggest that combining labeled stereo datasets with high-quality pseudo-stereo data can significantly enhance the generalization capability of stereo models. The diverse, complementary strengths of labeled and pseudo-stereo datasets allow the model to effectively learn detailed and generalized disparity features, ultimately leading to robust performance across various real-world scenarios.
6.5 Qualitative Results
In Figure 4, we present the qualitative results of Stereo Anything on five previously unseen datasets. Our model demonstrates strong robustness across various domains, including both indoor and outdoor scenes. Compared to the baseline model, NMRF-Stereo-SwinT [13], our approach consistently produces more accurate disparity maps, as indicated in the visualizations. This qualitative analysis underscores the model’s strong generalization capabilities and its ability to perform well across domains with substantial visual and environmental diversity, validating its effectiveness and potential for real-world stereo-matching applications.
7 Conclusion
In this paper, we introduce StereoAnything, a highly practical solution for robust stereo matching. We built a new synthetic dataset to better generalize across different scenarios and enhance the performance, called StereoCarla. Compared to existing datasets, it features unique view angles and baselines. Then, we investigated the effectiveness of labeled stereo datasets and pseudo stereo datasets generated using monocular depth estimation models to enhance stereo model generalization. Our experiments demonstrate that the quality and diversity of the datasets play a critical role in achieving robust stereo-matching performance across different domains. As a result, our StereoAnything achieves competitive performance across various benchmarks and real-world scenarios. These findings underscore the potential of hybrid training strategies that incorporate diverse data sources to enhance stereo model robustness, paving the way for future advancements in stereo matching.
Acknowledgements. This work was supported by the National Natural Science Foundation of China under Grant 62373356 and the Open Projects Program of the State Key Laboratory of Multimodal Artificial Intelligence Systems.
[…]
Supplementary Material
Figure 5: Disparity distribution across different datasets. We plot the percentage of pixels at different disparities over single datasets.
8 Disparity Distribute
Figure 5 plots the different disparity distributions in each existing labeled stereo dataset, highlighting the diversity in the range and frequency of disparities present. The figure presents the percentage of pixels at different disparity levels for well-known datasets.
9 More Implementation Details
Given the relatively low resolution of ImageNet21K, we resize the images while preserving their original aspect ratio to facilitate training on pseudo stereo datasets. Specifically, we ensure that the width is at least 768 pixels and the height is at least 384 pixels. This resizing strategy helps preserve the important visual features in the images while maintaining a resolution suitable for effectively training our model. The resolution of the input image is $352\times 640$.
10 More Details about StereoCarla
Table 11 shows the number of samples collected across different towns under various camera configurations. Figure 6 illustrates the left and right images at varying baselines, the corresponding disparity, and diagrams from various angles. These visualizations help in understanding how different baselines affect the stereo image pairs and the resulting disparity, providing insight into the geometry of the captured scenes.
town01 town02 town03 town04 town05 town06 town07 town10
normal 11700 12425 12235 12335 12420 12420 11090 12415
roll05 12425 12425 12330 9165 11470 12425 12430 12420
roll15 11750 12410 11475 12410 12415 12350 12380 12420
roll30 12195 12405 11165 12415 11930 12145 12420 9995
pitch00 12375 12410 5280 12375 12410 7450 9795 6490
pitch30 12415 12410 12410 11295 12400 12395 12425 6065
Table 11: StereoCarla samples distribution. Number of Samples Collected Across Different Towns and Camera Configurations.
11 Limitations and Future Works
Currently, our model is trained using over 2 million images. Although, in the current stage, this is sufficient to achieve unprecedented generalization performance, in the future, we plan to incorporate more unsupervised monocular data to train a more robust and effective model, further enhancing its performance and generalization capabilities.
12 More Qualitative Results
We present comprehensive qualitative results from five unseen test sets. Specifically, Figure 7 corresponds to KITTI 12, Figure 8 corresponds to KITTI 15, Figure 9 corresponds to Middlebury, Figure 10 corresponds to ETH3D, and Figure 11 corresponds to DrivingStereo. We compare our results with the state-of-the-art NMRFStereo model. Our method demonstrates higher disparity accuracy and greater robustness compared to NMRFStereo, indicating the effectiveness of our approach across diverse datasets.
Figure 6: Overall of StereoCarla datasets. The first row illustrates the left-eye image (1st column) and the right-eye image at varying baselines (2nd - 6th columns). The second row showcases the depth map (1st column) and corresponding disparity maps (2nd - 6th columns).
Figure 7: Qualitative results on KITTI12. NMRFStereo refers to NMRF-Stereo-SwinT.
Figure 8: Qualitative results on KITTI15. NMRFStereo refers to NMRF-Stereo-SwinT.
Figure 9: Qualitative results on Middlebburry. NMRFStereo refers to NMRF-Stereo-SwinT.
Figure 10: Qualitative results on ETH3D. NMRFStereo refers to NMRF-Stereo-SwinT.
Figure 11: Qualitative results on DrivingStereo. NMRFStereo refers to NMRF-Stereo-SwinT.

### v2

@@
 We extensively evaluate the zero-shot capabilities of our model on five public datasets, showcasing its impressive ability to generalize to new, unseen data (Figure 1).
-Code will be available at https://github.com/XiandaGuo/OpenStereo.
 Figure 1: StereoAnything demonstrates impressive generalization capabilities in different unseen scenes.
-11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding authors: long.chen@ia.ac.cn
+11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding authors
 1 Introduction
@@
 2.3 Multi-Dataset Training
-To tackle the challenge of generalization, some researchers have increasingly explored the use of multi-dataset training strategies.
+To tackle the challenge of generalization, some researchers have increasingly explored multi-dataset training strategies.
 By combining multiple datasets, models can be trained to capture diverse scene characteristics better, making them more robust.
@@
 Dataset Indoor Outdoor # Images
+LSUN [51] 9.8M
+ImageNet-21K [32] 13.1M
 BDD100K [52] 0.1M
+Google Landmarks [45] 4.97M
 Places365 [61] 2.1M
-Google Landmarks [45] 4.97M
-ImageNet-21K [32] 13.1M
 Table 2: Unlabeled monocular images. M refers to million.
@@
 The third rows depict left images from varied horizontal viewing angles and elevated viewpoints.
-To enhance the robustness of the stereo-matching model, we incorporate a large-scale collection of 20.27 million unlabeled monocular images sourced from four public datasets.
+To enhance the robustness of the stereo-matching model, we incorporate a large-scale collection of 30.07 million unlabeled monocular images sourced from four public datasets.
 A comprehensive overview of these datasets is provided in Table 2.
@@
 ETH3D [35] has 27/20 grayscale stereo image pairs for training and testing, collected from both indoor and outdoor scenes, annotated with semi-dense ground-truth labels obtained with a Faro Focus X 330 laser scanner.
-DrivingStereo [48] is a large-scale stereo dataset which includes over 180K images covering a diverse set of real-world driving scenarios.
+DrivingStereo [48] is a large-scale stereo dataset that includes over 180K images covering a diverse set of real-world driving scenarios.
 Sparse ground-truth annotations are obtained from LiDAR scans, being post-processed through a deep neural network.
@@
 LightStereo-L 6.41 6.40 17.51 11.33 21.74 12.68
-IGEV [46] 4.88 5.16 8.47 3.53 6.2 5.67
+IGEV [46] 4.88 5.16 8.47 3.53 6.20 5.67
 StereoBase [15] 4.85 5.35 9.76 3.12 11.84 6.98
@@
 Our training approach consistently delivers significant performance improvements across all datasets, showcasing its versatility and effectiveness.
-The most striking improvement is seen in the NMRFStereo-SwinT [13] variant shows particularly significant improvement, with the mean error reduced from 18.11 to 5.01.
+The most striking improvement is seen in the NMRFStereo-SwinT [13] variant, which shows particularly significant improvement, with the mean error reduced from 18.11 to 5.01.
 This notable gain is attributed to the increased data scale, which allows the SwinTransformer [26] backbone to unlock its full potential.
@@
 Table 7: Combinations of labeled stereo datasets.
+Figure 4: Qualitative results on five unseen datasets. NMRFStereo refers to NMRF-Stereo-SwinT [13].
 6.4.3 Training on pseudo stereo datasets.
-As shown in Table 8, we first validate the impact of using different monocular depth estimation models to generate pseudo stereo data on the BDD100k dataset.
+As shown in Table 8, we first validate the impact of using different monocular depth estimation models to generate pseudo stereo data on the BDD100k [52] dataset.
 Fine-tuning on the stereo datasets synthesized using MiDas [31], DepthAnytingV1 [49] (DAv1), and DepthAnytingV2 [50] (DAv2) leads to competitive results, demonstrating the effectiveness of using monocular depth models for pseudo stereo data generation.
 Specifically, the synthesized data from DAv2 [50] outperforms other monocular methods across all benchmarks.
-Figure 4: Qualitative results on five unseen datasets. NMRFStereo refers to NMRF-Stereo-SwinT [13].
 Then, we evaluate the impact of using different monocular datasets on the performance of the pseudo stereo data generation, based on the DAv2 [50] model.
 As shown in Table 9, the Google Landmarks [45] achieves the best performance overall, with a mean metric of 5.91.
-The slightly worse performance of ImageNet21K [32] (21K) compared to the GL [45] dataset can be attributed to the lower resolution of ImageNet21K images, which is smaller than our model’s input size.
-During the process, we resized the ImageNet21K [32] images to fit the input size of our model before running the monocular model to estimate depth, followed by generating the pseudo stereo pairs.
+The slightly worse performance of ImageNet21K [32] (21K) and LSUN [51] compared to the GL [45] dataset can be attributed to the lower resolution, which is smaller than our model’s input size.
+During the process, we resized the images of 21K [32] and LSUN [51] to fit the input size of our model before running the monocular model to estimate depth, followed by generating the pseudo stereo pairs.
 This resizing process might have led to a loss of fine details, thereby affecting the quality of the generated depth and resulting in slightly lower performance compared to GL [45].
@@
 SF [27] 8.67 7.46 16.36 23.46 34.58 18.11
+LSUN [51] 4.27 4.17 11.88 3.91 10.04 6.85
+21K [32] 4.10 4.43 10.44 4.25 9.08 6.46
 BDD [52] 3.69 3.87 11.75 3.92 9.55 6.56
 PL [61] 3.78 3.91 10.41 3.65 9.19 6.19
-21K [32] 4.10 4.43 10.44 4.25 9.08 6.46
 GL [45] 3.86 3.98 9.48 3.26 8.97 5.91